Hello guys! here is my question...
A car 15 feet long overtakes a truck 30 feet long which is traveling at the rate of 45 miles per hour. How fast must the car travel to pass the truck in 3 seconds?

Found 2 solutions by edjones, stanbon:
Answer by edjones(7311)

15+30=45' distance the car must travel to overtake the truck
It must be traveling 45/3=15'/sec faster than the truck.
45 mi/h=45*1.46666666667
=66'/sec
66+15
81'/sec or 55.2273 mph
.
Ed

Answer by stanbon(48558)

A car 15 feet long overtakes a truck 30 feet long which is traveling at the rate of 45 miles per hour. How fast must the car travel to pass the truck in 3 seconds?
--------------
In 3 seconds the truck advances (3/3600)*45 = 0.0375 miles or 198 feet.
The car must advance (198 + 30) = 228 ft in 3 seconds
228ft = 0.0431818.... miles
--------
EQUATION:
Rate = d/t = 0.0431818 miles / [3/3600 hr] = 51.82 mph
===============
Cheers,
Stan H.